If the hypotenuse of a triangle is 10 inches longer than the shorter side and 5 inches longer than the longer side.
What are the sides of the triangle?

Found 2 solutions by ankor@dixie-net.com, ikleyn:
Answer by ankor@dixie-net.com(22740)

If the hypotenuse of a triangle is 10 inches longer than the shorter side and 5 inches longer than the longer side.
What are the sides of the triangle?
:
using a^2 + b^2 = c^2
using the given information we can write these equation
a = c-10
b = c-5
then
(c-10)^2 + (c-5)^2 = c^2
FOIL
(c^2 - 20c + 100) + (c^2 - 10c + 25) = c^2
combine like terms
c^2 + c^2 - c^2 - 20c - 10c + 100 + 25 = 0
c^2 - 30c + 125 = 0
Can be factored to
(c-5)(c-25) = 0
Two solution but only
c = 25 will be valid, is the hypotenuse
and
a = 25 - 10
a = 15 is the shorter side
and
b = 25 - 5
b = 20 is the longer side

Answer by ikleyn(52778)

.

As the problem is worded, it makes me shudder.

Below I write how it should sound in normal Math language:

    +--------------------------------------------------------------------+
    |         If the hypotenuse of a triangle is 10 inches longer        |
    |   than the shorter leg and 5 inches longer than the longer leg,    |
    |             what are the sides of the triangle?                    |
    +--------------------------------------------------------------------+


My advise to the Math composer is to learn the basic terminology
of the subject and use it properly.